# TEncDM: Understanding the Properties of Diffusion Model in the Space of Language Model Encodings

### v1

@@ Abstract
Drawing inspiration from the success of diffusion models in various domains, numerous research papers proposed methods for adapting them to text data. Despite these efforts, none of them has managed to achieve the quality of the large language models. In this paper, we conduct a comprehensive analysis of key components of the text diffusion models and introduce a novel approach named Text Encoding Diffusion Model (TEncDM). Instead of the commonly used token embedding space, we train our model in the space of the language model encodings. Additionally, we propose to use a Transformer-based decoder that utilizes contextual information for text reconstruction. We also analyse self-conditioning and find that it increases the magnitude of the model outputs, allowing the reduction of the number of denoising steps at the inference stage. Evaluation of TEncDM on two downstream text generation tasks, QQP and XSum, demonstrates its superiority over existing non-autoregressive models.
TEncDM: Understanding the Properties of Diffusion Model
in the Space of Language Model Encodings
[…]
@@ 1 Introduction
Autoregressive (AR) large language models such as GPT-4 (OpenAI, 2023) or Llama 2 (Touvron et al., 2023) are the current gold standard in the text generation problem. They are capable of creating high-quality and coherent texts that are practically indistinguishable from the human ones. However, the disadvantage of this approach is the inability of the model to correct its own mistakes made during left-to-right generation. If a mistake occurs, it may spoil the subsequent text. In addition, the autoregressive method of token generation slows down the inference process as it requires performing a single model evaluation for each new token.
@@
Diffusion model is currently the state-of-the-art approach for data generation in image (Rombach et al., 2022; Betker et al., ), audio (Evans et al., 2024) and video (Blattmann et al., 2023) domains. They are a class of probabilistic generative models that are able to iteratively transfer noise to a representative sample of data. While some of the proposed text diffusion models are autoregressive Lovelace et al. (2022); Zhang et al. (2023), the majority of them are not and, by design, they have several advantages over AR language models. First, being non-autoregressive (NAR) models, they generate all the tokens simultaneously and can adjust any part of the sequence during the generation process. They also can be faster than AR models because the number of neural function evaluations for diffusion models depends on the number of denoising iterations rather than the length of the sequence. And given the possibility of distillation of diffusion models (Meng et al., 2023), the number of iterations can be greatly reduced.
@@
To date, a number of text diffusion models have been proposed, each based on substantially new ideas with little overlap with other methods. Some works replace Gaussian noise with categorical noise Hoogeboom et al. (2021); Austin et al. (2021), exploiting the discreteness of the text domain. Others train continuous diffusion on token embeddings Li et al. (2022); Lin et al. (2023); Gong et al. (2023) or text latent representations reduced in size Lovelace et al. (2022); Zhang et al. (2023). There are also differences in the way diffusion outputs are decoded back into text. Diffusion models trained on embeddings round their predictions to the nearest embeddings, while those that utilize small latent spaces decode the predictions with an AR model. This suggests that the scientific community has not found the most robust diffusion model design yet.
@@
In this paper, we attempt to better understand the specifics of text distribution models and identify best practices for their development. We investigate each component in detail: text encoding and decoding methods, diffusion model architecture, noise schedule, and self-conditioning (Chen et al., 2023). As a result, we combine all our findings in a method called Text Encoding Diffusion Model (TEncDM). It constructs the diffusion model, which operates in the latent space of the language model encodings (e.g. BERT (Devlin et al., 2019)). It also utilizes the Transformer-based (Vaswani et al., 2017a) decoder, which is able not only to decode the latents but also to improve the text quality. We do not use an AR decoder on purpose so as not to transfer the limitation of AR language models to the diffusion.
@@
We compare our approach with other works on two conditional text generation problems: paraphrasing and summarization, on which our method surpasses all on-autoregressive models. The main contributions of this work are as follows:
@@
  • We evaluate the importance of the decoder and conclude that its robustness to inaccuracies in the generated latents directly affects the generation quality. We then propose a decoder architecture and its training method that boosts the model performance.
  • We analyse in detail the effect of self-conditioning on the denoising process and show that self-conditioning increases the magnitude of model’s predictions, which in turn allows us to reduce the number of denoising steps during inference.
  • Through a thorough ablation study, we reveal that commonly used cosine and sqrt noise schedules do not introduce enough difficulty to the denoising task during training. We show that the addition of more noise significantly increases the model quality.
@@ 2 Problem Statement and Background
Autoregressive language models.
The most common approach for text generation is autoregressive (AR) left-to-right sampling of words. The idea is to approximate the factorised distribution $p(y)=\prod_{i=1}^{n}p(y_{i}|y_{<i})$ by learning a neural network $p_{\theta}(y_{i}|y_{<i})$. During generation, tokens are sampled sequentially with conditioning on the already generated ones.
@@
The standard diffusion models Ho et al. (2020); Song et al. (2021) learn to sample data from an unknown distribution by gradually denoising random Gaussian noise. The train procedure is defined through a forward diffusion process that satisfies $q(z_{t}|z_{0})=\mathcal{N}(\sqrt{\alpha_{t}}z_{0},(1-\alpha_{t})\mathbf{I})$, where $\alpha_{t}\in[0,1]$ is a predefined noise schedule, $t\in[0,1]$. The denoizing network (parameterized by $\theta$) is trained to reconstruct the original latent $z_{0}$ given the noisy latent $z_{t}$, as expressed in equation 1
$\mathcal{L}(\theta)=\mathop{\mathbb{E}}_{\varepsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim U[0;1]}[||z_{0}-\hat{z}_{\theta}(z_{t},t)||^{2}]$ (1)
Sampling procedure starts from a pure Gaussian noise $z_{T}\sim\mathcal{N}(0,\mathbf{I})$ and utilizes the denoizing network to iteratively generate latents $z_{t_{T-1}},...,z_{t_{1}}$, where $01=t_{T}>t_{T-1}>...>t_{1}=$ 0.
@@ 3 Related Work
The majority of proposed text diffusion models use embeddings of tokens to construct the continuous latent space (Li et al., 2022; Lin et al., 2023; Strudel et al., 2022; Gong et al., 2023; Wu et al., 2023). At the inference stage, to convert the latent predictions into text, they map each latent vector to a token corresponding to the nearest embedding.
Latent diffusion models.
Other studies suggest reducing the size of the latent space by training a diffusion model in the space of text autoencoder. PLANNER (Zhang et al., 2023) finetunes BERT (Devlin et al., 2019) to store all information in the first $k$ hidden state vectors from its final layer and use them as a latent space. LD4LG (Lovelace et al., 2022) trains the compression network to reduce both the length and dimensionality of the latent space sample. Both methods utilize autoregressive language models to decode the latents into text.
@@
Self-conditioning is a technique that significantly increases the performance of the diffusion model (Chen et al., 2023; Strudel et al., 2022; Lovelace et al., 2022). Usually the model is conditioned only on the latent variable $z_{t}$ and the current timestep $t$ as $\hat{z}_{0}^{t}=\hat{z}_{\theta}(z_{t},t)$. Self-conditioning proposes to also condition the model on the estimation of data sample from the previous timestep during generation in order to improve the prediction at the current timestep, $\hat{z}_{0}^{t}=\hat{z}_{\theta}(z_{t},t,\hat{z}_{0}^{t-1})$.
Although widely used, no analysis has been conducted to determine why this method is effective or how it impacts the generation process.
@@
Noise scheduler is a key component of a diffusion model that controls the amount of noise added on each timestep. Previous research (Li et al., 2022; Gao et al., 2023; Ye et al., 2023) has highlighted that the standard noise schedulers used for image diffusion models are unsuitable for the textual domain. Due to the discrete nature of the texts, it is unlikely that an addition of a small amount of noise to a latent will change its nearest text in the latent space. Therefore, to increase the difficulty of the denoising task for the model, the mentioned works recommend adding more noise on iterations that are close to 0.
@@ 4 Understanding Text Diffusion
Most diffusion models utilize token embeddings to map text into a continuous latent space. However, this approach is not optimal because the embeddings do not convey contextual information. This requires the diffusion model to independently search for it to retrieve ambiguous tokens. To simplify the task, instead of embeddings, we can use the final layer outputs of a pre-trained language model (e.g. BERT). They contain this information and, thus, should be more suitable for training the diffusion model. We refer to these outputs as encodings.
Experimental results confirming our intuition are presented in Section 7.3. It is worth noting that the use of encodings does not slow down the generation process, as we need to compute them only during the training.
To improve the quality even further, it is possible to fine-tune the encoder, but we choose not to in order to avoid overcomplicating the approach. Investigation into fine-tuning is left for the future work.
@@
Following the recommendations of previous works (Li et al., 2022; Wu et al., 2023; Ye et al., 2023), we used sqrt noise scheduler that increases the amount of noise added to the diffusion model inputs during training beyond the amount of typically used cosine noise scheduler (Han et al., 2022; Lovelace et al., 2022; Strudel et al., 2022; Zhang et al., 2023). However, our experiments led us to conclusion that encoding-based diffusion model requires even more noise for successful training. We hypothesize that this is due to the presence of contextual information in the encodings, which simplifies the denoising task.
@@ 5 Methodology
The design of TEncDM is depicted on Figure 1. It consists of three parts – diffusion encoder $E_{diff}$, diffusion model $\hat{z}_{\theta}$ and decoder $D$. For the conditional generation, we also add conditional encoder $E_{cond}$, which encodes an input text. Its output is provided to the diffusion model and decoder through cross-attention.
@@ 5.1 Diffusion encoder, $E_{diff}$
We use pre-trained Transformer-based Vaswani et al. (2017a) language model $E_{diff}$, which we call diffusion encoder, to encode text $y$ into the latent space $z$. Encoding of text does not change the length of the sequence. In order to align all texts in length, we add paddings to the end of short texts. After encoding the text, the encodings of all special tokens are replaced by their corresponding embeddings. This is necessary because diffusion model does not use an attention mask during training, which means that the reconstruction loss is calculated for both text and special tokens. However, special token encodings usually contain meaningless values, because encoder does not learn to store useful information in them. Therefore, minimization of reconstruction loss for these encodings only harms the diffusion training process. Embeddings of special tokens, on the other hand, only contain information about the token itself and the diffusion model recovers them much easier.
@@ 5.2 Decoder, $D$
The decoder $D$ is required to convert latent variables generated by diffusion model into textual output. Although a basic linear decoder can effectively reconstruct tokens with high accuracy, we employ the BERT Devlin et al. (2019) architecture for the decoder to provide it with the ability to capture context information and rectify potential errors originating from the diffusion model.
@@ 5.3 Diffusion model, $\hat{z}_{\theta}$
We train the diffusion model using the variance preserving scheme, discussed in Song et al. (2021). To achieve zero mean and unit variance we normalize the latent variables $z_{0}$ coordinate-wise, using the statistics from the training set.
@@
We adopt the noise scheduler from (Hoogeboom et al., 2023) and use the following equation for $\alpha_{t}$:
@@
Following the previous approaches (Lovelace et al., 2022; Strudel et al., 2022) we incorporate self-conditioning into the diffusion model. In order to make the model utilize the data sample estimation from the previous generation step, we modify the training procedure.
According to Chen et al. (2023) we design the training process to emulate the inference behavior. On each training iteration with the probability $p=0.5$ the prediction is computed with the self-conditioning set to zero $\bar{z}_{0}^{t}=z_{\theta}(z_{t},t,0)$. And, with probability $(1-p)=0.5$ we first calculate $\bar{z}_{0}^{t}=z_{\theta}(z_{t},t,0)$ and then use it as an estimation of the data sample to obtain a second prediction $\tilde{z}_{0}^{t}=z_{\theta}(z_{t},t,\text{SG}(\bar{z}_{0}^{t}))$, where SG is the stop-gradient function that does not allow the gradient to flow through $\bar{z}_{0}^{t}$. The diffusion model is optimized using the output $\bar{z}_{0}^{t}$ in the former scenario and $\tilde{z}_{0}^{t}$ in the latter. This training strategy allows the model to accurately approximate $z_{0}$ both with and without self-conditioning. We implement self-conditioning in a same manner as conditioning on timestep. For each diffusion model layer we pass the data estimation through a single linear layer and add it to the hidden state vectors.
@@ 5.4 Generation process
5.5 Conditional generation
For the conditional generation we keep the framework design similar to unconditional generation. The only difference is that we add conditional encoder to process the input text and provide both diffusion model and decoder with its output via cross-attention. Implementation details can be found in Appendix E.

### v2

@@ Abstract
This paper presents the Text Encoding Diffusion Model (TEncDM), a novel approach to diffusion modeling that operates in the space of pre-trained language model encodings. In contrast to traditionally used embeddings, encodings integrate contextual information. In our approach, we also employ a transformer-based decoder, specifically designed to incorporate context in the token prediction process. We conduct a comprehensive examination of the influence of the encoder, decoder, noise scheduler, and self-conditioning on zero-shot generation. Furthermore, we compare TEncDM with previous approaches on three conditional text generation tasks: QQP, XSum, and Wiki-Auto. The results show that TEncDM exhibits superior performance compared to existing non-autoregressive diffusion models.
@@ 1 Introduction
Autoregressive (AR) large language models such as GPT-4 (OpenAI 2023) or Llama 3 (Dubey et al. 2024) are the current gold standard in the text generation problem. They are capable of creating high-quality and coherent texts that are practically indistinguishable from the human ones. However, the disadvantage of this approach is the inability of the model to correct its own mistakes made during left-to-right generation. If a mistake occurs, it may spoil the subsequent text. In addition, the autoregressive method of token generation slows down the inference process as it requires performing a single model evaluation for each new token.
@@
Diffusion modeling is currently the state-of-the-art approach for data generation in image (Rombach et al. 2022; Podell et al. 2023), audio (Evans et al. 2024) and video (Blattmann et al. 2023) domains. Researchers now attempt to adapt it also for text generation (Li et al. 2022; Gong et al. 2023; Karimi Mahabadi et al. 2024). Diffusion models are a class of probabilistic generative models that are able to iteratively transfer noise to a representative sample of data. While some of the proposed text diffusion models are autoregressive (Lovelace et al. 2023; Zhang et al. 2023), the majority of them are not and, by design, they have several advantages over AR language models. First, being non-autoregressive (NAR) models, they generate all the tokens simultaneously and can adjust any part of the sequence during the generation process. They also can be faster than AR models because the number of neural function evaluations for diffusion models depends on the number of denoising iterations rather than the length of the sequence. And given the possibility of distillation of diffusion models (Meng et al. 2023), the number of iterations can be greatly reduced.
@@
To date, a number of text diffusion models have been proposed, each based on substantially new ideas with little overlap with other methods. Some works replace Gaussian noise with categorical noise (Hoogeboom et al. 2021; Austin et al. 2021), exploiting the discreteness of the text domain. Others train continuous diffusion on token embeddings (Li et al. 2022; Lin et al. 2023; Gong et al. 2023), on text latent representations reduced in size (Lovelace et al. 2023; Zhang et al. 2023) or on the probability simplex (Karimi Mahabadi et al. 2024; Han, Kumar, and Tsvetkov 2023). There are also differences in the way diffusion outputs are decoded back into text. Diffusion models trained on embeddings round their predictions to the nearest embeddings, while those that utilize small latent spaces decode the predictions with an AR model. This suggests that the scientific community has not yet found the most robust design of a diffusion model for text.
@@
In this paper, we attempt to better understand the specifics of continuous text diffusion models trained in the latent space of token embeddings and identify best practices for their development. We argue that the use of raw embeddings as latents is suboptimal, and that this approach can be enhanced by first extracting context information from embeddings with a language model encoder and training a diffusion model in this latent space. We also investigate several diffusion components in detail: text decoding methods, diffusion model architecture, noise schedule, and self-conditioning (Chen, Zhang, and Hinton 2023). As a result, we combine all our findings in a method called Text Encoding Diffusion Model (TEncDM).
@@
We compare our approach with other non-autoregressive diffusion models trained in the space of embeddings or encodings on three conditional text generation problems – paraphrasing, summarization, and text simplification – and show its superiority over other methods. The main contributions of this work are as follows:
@@
  • We evaluate the importance of the text decoder and conclude that its robustness to inaccuracies in the generated latents directly affects the generation quality. We propose a Transformer-based decoder and its training method that boosts the model performance.
  • We analyse in detail the effect of self-conditioning and noise schedules on the denoising process and show their effect on the model quality.
@@ 2 Problem Statement and Background
The standard diffusion models (Ho, Jain, and Abbeel 2020; Song et al. 2021) learn to sample data from an unknown distribution by gradually denoising random Gaussian noise. The training procedure is defined through a forward diffusion process that satisfies $q(z_{t}|z_{0})=\mathcal{N}(\sqrt{\alpha_{t}}z_{0},(1-\alpha_{t})\mathbf{I})$, where $\alpha_{t}\in[0,1]$ is a predefined noise schedule, $t\in[0,1]$ and $\alpha_{t}>\alpha_{t+\Delta t}$. The denoising network (parameterized by $\theta$) is trained to reconstruct the original latent $z_{0}$ given the noisy latent $z_{t}$, as expressed in equation 1.
$\mathcal{L}(\theta)=\mathop{\mathbb{E}}_{\varepsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim U[0;1]}[\|z_{0}-\hat{z}_{\theta}(z_{t},t)\|^{2}]$ (1)
Sampling procedure starts from a pure Gaussian noise $z_{T}\sim\mathcal{N}(0,\mathbf{I})$ and utilizes the denoising network to iteratively generate latents $z_{t_{T-1}},...,z_{t_{1}}$, where $01=t_{T}>t_{T-1}>...>t_{1}=$ 0.
@@ 3 Related Work
The majority of proposed text diffusion models use embeddings of tokens to construct the continuous latent space (Li et al. 2022; Lin et al. 2023; Strudel et al. 2022; Gong et al. 2023; Wu et al. 2023). At the inference stage, to convert the latent predictions into text, they map each latent vector to a token corresponding to the nearest embedding.
@@
Self-conditioning is a technique that significantly increases the performance of the text diffusion model (Chen, Zhang, and Hinton 2023; Strudel et al. 2022; Lovelace et al. 2023). Usually the model is conditioned only on the latent variable $z_{t}$ and the current timestep $t$ as $\hat{z}_{0}^{t}=\hat{z}_{\theta}(z_{t},t)$. Self-conditioning proposes to also condition the model on the estimation of data sample from the previous timestep during generation in order to improve the prediction at the current timestep, $\hat{z}_{0}^{t}=\hat{z}_{\theta}(z_{t},t,\hat{z}_{0}^{t-1})$.
Although widely used, no analysis has been conducted to determine why this method is effective or how the generation process is altered by its application.
@@
Noise scheduler is a key component of a diffusion model that controls the amount of noise added on each timestep. Previous research (Li et al. 2022; Gao et al. 2024; Ye et al. 2024) has highlighted that the standard noise schedulers used for image diffusion models are unsuitable for the textual domain. Due to the discrete nature of the texts, it is unlikely that an addition of a small amount of noise to a latent will change its nearest text in the latent space. Therefore, to increase the difficulty of the denoising task for the model, the mentioned works recommend adding more noise on iterations that are close to 0.
@@ 4 Understanding Text Diffusion
Most diffusion models utilize token embeddings to map text into a continuous latent space. However, this approach is not optimal because the embeddings do not convey contextual information. This requires the diffusion model to independently search for it to retrieve ambiguous tokens. To simplify the task, instead of embeddings, we can use the final layer outputs of a pre-trained language model (e.g. BERT). They contain contextual information and, thus, should be more suitable for training the diffusion model. We refer to these outputs as encodings.
Experimental results confirming our intuition are presented in Section 7.2. It is worth noting that the use of encodings does not slow down the generation process, as we need to compute them only during the training. To improve the quality even further, it is possible to fine-tune the encoder, but we choose not to in order to avoid overcomplicating the approach. Investigation into fine-tuning is left for the future work.
@@
Following the recommendations of previous works (Li et al. 2022; Wu et al. 2023; Ye et al. 2024), we used sqrt noise scheduler that increases the amount of noise added to the diffusion model inputs during training beyond the amount of typically used cosine noise scheduler (Han, Kumar, and Tsvetkov 2023; Lovelace et al. 2023; Strudel et al. 2022; Zhang et al. 2023). However, our experiments led us to conclusion that encoding-based diffusion model requires even more noise for successful training. We hypothesize that this is due to the presence of contextual information in the encodings, which simplifies the denoising task.
@@ 5 Methodology
The design of TEncDM is depicted on Figure 1. It consists of three parts – diffusion encoder $E_{diff}$, diffusion model $\hat{z}_{\theta}$ and decoder $D$. For the conditional generation, we also add condition encoder $E_{cond}$, which encodes an input text. Its output is provided to the diffusion model and decoder through cross-attention.
@@ 5.1 Diffusion encoder, $E_{diff}$
We use pre-trained Transformer-based (Vaswani et al. 2017b) language model $E_{diff}$, which we call diffusion encoder, to encode text $y$ into the latent variable $z$. Encoding of text does not change the length of the sequence. In order to align all texts in length, we add paddings to the end of short texts. After encoding the text, the encodings of all special tokens are replaced by their corresponding embeddings and padding encodings are replaced with zeros. This is necessary because diffusion model does not use an attention mask during training, which means that the reconstruction loss is calculated for both text and special tokens. However, special token encodings usually contain meaningless for diffusion model values. Therefore, minimization of reconstruction loss for these encodings only harms the training process. Embeddings of special tokens, on the other hand, only contain information about the token itself and the diffusion model recovers them much easier. During training we do not update the weights of the encoder in order to keep the approach simple.
@@ 5.2 Decoder, $D$
The decoder $D$ is required to convert latent variables generated by diffusion model into textual output. Although a basic linear decoder can effectively reconstruct tokens with high accuracy, we employ the BERT-type (Devlin et al. 2019) architecture for the decoder to provide it with the ability to capture context information and rectify potential mistakes originating from the diffusion model. Note that we do not use an AR decoder on purpose so as not to transfer the limitations of AR language models to the diffusion model.
@@ 5.3 Diffusion model, $\hat{z}_{\theta}$
We train the diffusion model using the variance preserving scheme, discussed in (Song et al. 2021). To achieve zero mean and unit variance we normalize the latent variables $z_{0}$ coordinate-wise, using the statistics from the training set.
@@
We adopt the noise scheduler from (Hoogeboom, Heek, and Salimans 2023) and use the following equation for $\alpha_{t}$:
@@
Following the previous approaches (Lovelace et al. 2023; Strudel et al. 2022) we incorporate self-conditioning into the diffusion model. In order to make the model utilize the data sample estimation from the previous generation step, we modify the training procedure.
According to (Chen, Zhang, and Hinton 2023) we design the training process to emulate the inference behavior. On each training iteration with the probability $p=0.5$ the prediction is computed with the self-conditioning set to zero $\bar{z}_{0}^{t}=z_{\theta}(z_{t},t,0)$. And, with probability $(1-p)=0.5$ we first calculate $\bar{z}_{0}^{t}=z_{\theta}(z_{t},t,0)$ and then use it as an estimation of the data sample to obtain a second prediction $\tilde{z}_{0}^{t}=z_{\theta}(z_{t},t,\text{SG}(\bar{z}_{0}^{t}))$, where SG is the stop-gradient function that does not allow the gradient to flow through $\bar{z}_{0}^{t}$. The diffusion model is optimized using the output $\bar{z}_{0}^{t}$ in the former scenario and $\tilde{z}_{0}^{t}$ in the latter. This training strategy allows the model to accurately approximate $z_{0}$ both with and without self-conditioning. We implement self-conditioning in a same manner as conditioning on timestep. For each diffusion model layer we pass the data estimation through a single linear layer and add it to the hidden state vectors.
@@ 5.4 Generation process
For the conditional generation we keep the framework design similar to unconditional generation. The only difference is that we add condition encoder to process the input text and provide both diffusion model and decoder with its output via cross-attention. Implementation details can be found in Appendix E.
Encoder ppl $\downarrow$ mem $\downarrow$ div $\uparrow$ mauve $\uparrow$
ROCStories
BERT emb $48.9_{.36}$$371_{.003}.$$324_{.002}.$$600_{.016}.$
BERT $29.1_{.89}$${.453}_{.003}.$${.295}_{.002}.$$\textbf{.762}_{.043}.$
RoBERTa $\textbf{28.3}_{.33}$${.443}_{.003}.$${.302}_{.002}.$${.647}_{.019}.$
T5 ${31.3}_{.54}$$427_{.003}.$$312_{.004}.$$706_{.024}.$
BART $34.1_{.52}$$441_{.006}.$$299_{.005}.$$705_{.030}.$
Source text $21.7$$365$$403$876.876.876.876
Wikipedia
BERT emb $156.1_{1.8}$$263_{.004}.$$517_{.002}.$$378_{.055}.$
BERT $\textbf{104.4}_{2.1}$$286_{.002}.$$504_{.003}.$$\textbf{.874}_{.011}.$
Source text $37.3$$122$$615$957.957.957.957
Table 1: Comparison of diffusion encoders.